OpenStudy (elifbestegul):

Find x. -> 3e^x - 5 = -2e^-x

OpenStudy (irishboy123):

make the sub \( u = e^x\) so \( e^{-x} = 1/u\) and solve it as a quadratic in u then unwind the sub back into x

OpenStudy (elifbestegul):

okay I am trying in a few seconds @irishboy123

OpenStudy (elifbestegul):

I have done it thanks :) @irishboy123